3. There are 20 black cell phones and 30 white cell phones in a store. An employee t phones at random. Find the probability that (a) there will be exactly 4 black cell phones among the chosen phones; (b) there will be less than 3 black cell phones among the chosen phones.
Answer #1

Total number of phone = 20 + 30-50 Number of way of choosing 10 phone out of 50 phone -50 C1o (a) Number of way of choosing exactly 4 black cell phone- C4×30 C6 Hence probability - (b) Nurnber of way of choosing less than 3 black cell phone- Cox 0 C10+20 G Hence probability = 20C4x30 C6 0.28 50C10 C9+20 C2 x30 Cs 20C0 ×30 C10 +20 Cl ×30C9 +20 C2 ×30 Cs 50 C = 0.1390

> Answer in b) is correct but should say 50 C 10 in the numerator

Steinar Bjornsson Thu, Jan 13, 2022 3:10 AM